3. Comparing salaries from different times Consider golfers who led the Professional Golfers' Association of America (PGA) in winnings at different points in time. Note that the winnings are nominal figures (unadjusted for inflation). 


To convert the original earnings of Palmer, Trevino, and Norman, use the formula for converting dollar figures from an earlier era into year 2016 U.S. dollars. Using those figures, fill in the following table, making sure to round your responses to the nearest U.S. dollar. 

Nominal Winnings U.S. CPI Year Golfer (Dollars) (1983 100) Winnings in 2016 Dollars 1963 Arnold Palmer 128,230 30.6 1970 Lee Trevino 157,037 38.8 Greg Norman 1986 653,296 109.6 2016 9,365,185 Dustin Johnson 240.0 9,365,185 


True or False: According to the previous table, the golfer with the highest PGA winnings in nominal dollars is not the same as the golfer with the highest PGA winnings after adjusting for inflation. True False
Answer #1

Winnings in 2016 dollars = Nominal winnings*(CPI in 2016/CPI in current year)

1963: 128,230*(240/30.6) = 1,005,726

1970: 157,037*(240/38.8) = 971,363

1986: 653,296*(240/109.6) = 1,430,575

False
(Dustin Johnson is the golfer with highest PGA winnings in both.)